OpenStudy (anonymous):

We toss two dice and record the sum of the numbers on the two dice. Then we repeat this 1,000 times. How many times do we expect to have the sum of the two dice equal to 4?

OpenStudy (anonymous):

it will be equal 4 when we have: 1_3 , 3_1, 2_2, 2_2 so the probability that the sum will be 4 is: 4/36=1/9 so if you repeat it 1000 times: 1000*1/9=1000/9